I need to find the slope if it exists of the line containing the pair of points
(-14.3,15.7) and (-11.8,2.5)
Thank you

Answer by haileytucki(390)

(-14.3,15.7),(-11.8,2.5)
Slope is equal to the change in y over the change in x, or 'rise over run'.
m=(change in y)/(change in x)
The change in x is equal to the difference in x-coordinates (also called run), and the change in y is equal to the difference in y-coordinates (also called rise).
m=(y2-y1)/(x2-x1)
Substitute in the values of x and y into the equation to find the slope.
m=(2.5-(15.7))/(-11.8-(-14.3))
Multiply -1 by each term inside the parentheses.
m=(2.5-(15.7))/(-11.8+14.3)
Add 14.3 to -11.8 to get 2.5.
m=(2.5-(15.7))/(2.5)
Reduce the expression ((2.5-(15.7)))/(2.5) by removing a factor of from the numerator and denominator.
m=0.4(2.5-(15.7))
Multiply -1 by the 15.7 inside the parentheses.
m=0.4(2.5-15.7)
Subtract 15.7 from 2.5 to get -13.2.
m=0.4(-13.2)
Multiply 0.4 by each term inside the parentheses.
m=-5.28